OpenStudy (anonymous):

Write an equation in slope-intercept form for an equation of a line perpendicular to y = -3x +1 and that intersects at point (1, 2/3)

OpenStudy (anonymous):

y = 1/3x +1 y + x / -1/3 =1

OpenStudy (anonymous):

sry y + x/-3 =1

OpenStudy (anonymous):

fantastic thnx!

OpenStudy (anonymous):

Ishaan- I got the first part finding the perpendicular slope, but what did you do in the second?

OpenStudy (anonymous):

i think the answer is y=1/3x-1 is that true?

OpenStudy (anonymous):

see , i did tht wrong again y = mx +c perpendicular line should be y = -1/mx +b where b is const now when u obtain all the other parts except b then wt u do is pt the pts in the eq and get the value of b in my case y - x/3 = b so pts frm which line passes (1 , 2/3) 2/3 - 1/3 =b 1/3=b y - x/3 = 1/3 sry i did like 4 times this question ..and just messed it up sry